OpenStudy (firejay5):

Solve the absolute. There should be 2, if not 1. Show work and instructions: 3|2A + 7| = 3A + 12

hero (hero):

1. Divide both sides by 3. Let me know what you get

OpenStudy (firejay5):

2A + 7 = 3A + 4

hero (hero):

You divided both sides by 3. That doesn't mean get rid of the absolute value just yet.

OpenStudy (firejay5):

oh okay

hero (hero):

So what do you have so far?

OpenStudy (firejay5):

|2A + 7| = 3A + 4

hero (hero):

Okay, so for the next step, apply this rule: If |a + b| = c , then a + b = ±c

OpenStudy (firejay5):

I just need the steps to the problem

hero (hero):

And I am guiding you through the steps.

OpenStudy (firejay5):

yes, but a + b = c will confuse me

hero (hero):

I wonder what's so difficult about applying the rule so that afterwards you get 2A + 7 = ±(3A + 4)

OpenStudy (firejay5):

I just need the steps and answer or I won't get it. Yes I got answer

OpenStudy (firejay5):

I am in algebra 2, we didn't have to put +/- by anything unless it's positive/negative

OpenStudy (anonymous):

so the end is 2A + 7 = 3A + 4 and 2A + 7 = -3A - 4 just solve those equations for A

OpenStudy (firejay5):

which A do I +/-

OpenStudy (anonymous):

Doesn't matter. As long as you put the A's on one side and the constants on the other (the numbers without variables) So A = 3 and 2A + 3A = -4 - 7 5A = -11 A = -11/5

OpenStudy (firejay5):

can you work out the problem, so I could see

OpenStudy (firejay5):

I got A = -3

OpenStudy (anonymous):

2A - 3A = 4 - 7 -A = -3 A = 3

OpenStudy (firejay5):

I am getting a positive A not a negative

OpenStudy (anonymous):

2A - 3A = -A or are you not getting 2A - 3A

OpenStudy (firejay5):

no I am it's just I missed interperted

OpenStudy (firejay5):

it

OpenStudy (firejay5):

sorry

OpenStudy (anonymous):

so do you know how I got -A now?

OpenStudy (firejay5):

yea! I am sorry @ArkGoLucky

OpenStudy (anonymous):

no no it's fine

OpenStudy (anonymous):

absolute values can be confusing

OpenStudy (firejay5):

I just calculated it wrong cause I am tired